A check-processing center uses exponential smoothing to forecast the number of incoming checks each month. The number of checks received in June was

38 million, while the forecast was 43 million. A smoothing constant of 0.25 is used.

a) Using exponential smoothing and given

alphaα , the forecast for the month of July = ?

million checks received (round your response to one decimal place).
Answer #1

Using the exponential smoothing method the formula to calculate the forecast is as follows :

Ft = F(t-1) + [A(t-1) - F(t-1)]

Where Ft = forecast for period t

A(t-1)= actual value for period previous to t

F(t - 1)= forecast for period previous to t

= smoothing constant

So here,

A(t-1) = actual number of check received in June = 38 million

F(t-1) = forecast for June = 43 million

= smoothing constant = 0.25

Forecast for July = 43 + 0.25(38-43)

= 43+[0.25 × (-5)]

= 43 +(-1.25)

= 43-1.25

= 41.8 millions